# PromptReps: Prompting Large Language Models to Generate Dense and Sparse Representations for Zero-Shot Document Retrieval

### v1

@@ 4 Experimental setup
Dataset and evaluation: We evaluate the document ranking effectiveness of both baseline methods and our proposed PromptReps using the BEIR benchmark Thakur et al. (2021). This benchmark encompasses various IR tasks, providing a heterogeneous evaluation environment. We report nDCG@10 scores, which is the official evaluation metric employed by BEIR.
Baselines: We compare PromptReps with strong unsupervised first-stage retrievers including BM25, a classic term frequency-based sparse retrieval method, and E5-PTlarge Wang et al. (2024a), a state-of-the-art BERT-based Devlin et al. (2019) dense embedding method trained on 1.3 billion carefully crafted unsupervised text pairs. LLM2Vec BehnamGhader et al. (2024), an LLM-based dense embedding method trained with bi-directional attention, masked next token prediction, and SimCSE Gao et al. (2021) on the Wikipedia corpus.
Implementation of PromptReps: PromptReps is implemented using three base LLMs: Mistral-7b-Instruct-v0.2444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), Llama3-8B-Instruct [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, Phi-3-mini-4k-instruct [^6] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024), and Llama3-70B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct AI@Meta (2024). Dense and sparse document and query encoding are implemented using the Huggingface Transformers library Wolf et al. (2020) and the Tevatron toolkit Gao et al. (2023b). Faiss library Douze et al. (2024) is used to build the ANN index with cosine similarity as the embedding distance metric, and Pyserini Lin et al. (2021) is utilized to construct the inverted index for sparse retrieval. For the dense and sparse ranking hybrid, the Ranx library Bassani and Romelli (2022) is employed. In our experiments, we report dense only, sparse only, and the full hybrid results. In addition, we also report the results of further interpolating BM25 scores with hybrid scores for PromptReps based on Llama3-70B-Instruct.
@@ 5 Results
We present our results in Table 1. The first observation highlights that BM25 is a very strong zero-shot retrieval method, capable of outperforming LLM2Vec, based on the Mistral-7B LLM, across numerous datasets, achieving a higher average nDCG@10 score. This outcome implies that even with a large-size LLM, bi-directional attention enabled, additional pre-training, and SimCSE-based unsupervised contrastive training, there remains a gap in transforming a decoder-only LLM into a robust retrieval method.
@@
For our proposed PromptReps, when using Mistral-7B-Instruct LLM, the dense-only retriever performs poorly. However, although still worse than LLM2Vec, the sparse-only retriever performs much better than dense. This finding suggests that, with Mistral-7B-Instruct as the base LLM, using the logits generated by our prompt produces better document representations than hidden states for document retrieval. The hybrid retrieval performs worse than sparse-only in this case due to the poor performance of dense retrieval.
Surprisingly, when implementing PromptReps with Phi-3-mini-4k-instruct achieved much higher dense-only retrieval effectiveness, closing to its sparse retrieval effectiveness and leading to an impressive hybrid retrieval average score that surpassed LLM2Vec based on Mistral-7B, even though it is only half the size.
@@
Interestingly, when implementing PromptReps with the state-of-the-art Llama3-8B-Instruct LLM, the dense-only retrieval further improved and achieved a higher score than that of sparse. The sparse-only retrieval effectiveness also improved, but the improvement was much more marginal compared to the improvement on dense. The hybrid of dense and sparse brings notable retrieval effectiveness improvement, surpassing BM25 and approaching the state-of-the-art E5-PTlarge. Notably, this is achieved without any form of extra training but solely relying on prompts.
@@
The scaling law of LLM Kaplan et al. (2020) is also applied here. When changing Llama3-8B-Instruction to Llama3-70B-Instruction, the retrieval effectiveness of our PromptReps further improved, with the dense-only approach surpassing LLM2Vec and the hybrid approach surpassing E5-PTlarge. When further interpolating the hybrid ranking with BM25 ranking, the average nDCG@10 achieved a remarkable score of $50.0750.07$. Our results demonstrate that it is possible to build a very strong hybrid retrieval system without the need for any unsupervised or supervised training.
A note on query latency. The best performance of our PromptReps is achieved in the hybrid retrieval setting. Compared to previous works which are dense retrieval only, our method seems to have a higher query latency. However, PromptReps actually only requires a little query latency overhead if implementing dense and sparse search in parallel. The cost of obtaining both dense and sparse representations of our method only requires a single LLM forward inference; the only extra computation is the dot product of the dense vector with the token embeddings, which is very fast on GPU. For document search, since we heavily sparsified the sparse representation, in our experiments, our sparse retrieval is much faster than BM25, and the bottleneck is the dense retrieval. Since the dense and sparse search could be run in parallel and the hybrid operation is just a linear interpolation of both rankings (very fast on CPU), the query latency will depend on the dense retrieval latency, thus very close to the previous methods.
6 Discussion
Our work demonstrates that prompt engineering for generative LLMs is not only a powerful method for natural language generation tasks but also capable of generating robust dense (embedding) and sparse (bag of words) representations for input texts. We believe this important finding could lead to a series of new research directions.
For example, techniques like few-shot in-context learning Brown et al. (2020), chain-of-thought prompting Wei et al. (2022), and auto-prompt optimization methods Yang et al. (2024); Fernando et al. (2023), which have proven to be very effective in text-generation tasks, could potentially be leveraged here to enhance embedding generation.
Moreover, it has been shown that the instruction-following ability of LLMs could be transferred to embedding models with synthetic instruction fine-tuning data Wang et al. (2024b). In our work, we always keep the instruction prompt consistent across different information retrieval tasks, which could be sub-optimal. It would be very interesting to investigate how to customize instructions for PromptReps to generate embeddings specific to different domains, tasks, or even to multi-lingual and cross-lingual retrieval settings.
Finally, our prompting method could be seen as a simple approach to obtaining a better initialization of LLM-based embedding models, which is much more cost-effective than methods requiring further pre-training BehnamGhader et al. (2024); Li et al. (2023). All the previous contrastive pre-training with paired text data and synthetically generated data could be applied on top of our method and could potentially yield improved LLM-based embedding models. We leave these aspects for future work.
@@ 7 Conclusion
In this paper, we introduced PromptReps, a simple yet effective method that prompts LLMs to generate dense and sparse representations for zero-shot document retrieval without any further unsupervised or supervised training. Our work reveals that modern LLMs are effective text encoders by themselves, and prompt engineering is sufficient to stimulate their text encoding ability.

### v2

@@ 4 Experimental setup
Dataset and evaluation: We evaluate the document ranking effectiveness of both baseline methods and our proposed PromptReps using MSMARCO Nguyen et al. (2016) passage retrieval, TREC deep learning Craswell et al. (2020) and BEIR Thakur et al. (2021). These datasets encompass various IR tasks, providing a heterogeneous evaluation environment. For MSMARCO we report MRR@10 and for TREC deep learning and BIER we report nDCG@10 scores, the commonly employed evaluation measures for these datasets.
Baselines: We compare PromptReps with strong unsupervised first-stage retrievers including BM25, a classic term frequency-based sparse retrieval method, and E5-PTlarge Wang et al. (2024a), a state-of-the-art BERT-based dense embedding method trained on 1.3B carefully crafted unsupervised text pairs. LLM2Vec BehnamGhader et al. (2024), a Llama3-8B-Instruct LLM-based dense embedding method trained with bi-directional attention, masked next token prediction, and SimCSE Gao et al. (2021) on the Wikipedia corpus.
Implementation of PromptReps: PromptReps is implemented using four base LLMs: Mistral-7b-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), Phi-3-mini-4k-instruct [^7] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024), Llama3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, and Llama3-70B-Instruct [^9] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct AI@Meta (2024). Dense and sparse document and query encodings are implemented using the Huggingface Transformers library Wolf et al. (2020) and the Tevatron toolkit Gao et al. (2023b). The Faiss library Douze et al. (2024) is used to build the ANN index with cosine similarity as the embedding distance metric, and Pyserini Lin et al. (2021) is utilized to construct the inverted index for sparse retrieval. For the dense and sparse ranking hybrid, the Ranx library Bassani and Romelli (2022) is employed. In our experiments, we report dense only, sparse only, and the full hybrid results.
@@ 5 Results
We start by showing our overall results on the BEIR dataset, which we treated as test set; we then analyse choices in instantiation of PromptReps, including different variations in the prompt using the MS MARCO and TREC deep learning datasets, which we used as development datasets to inform the choices we made to run PromptReps on BEIR.
5.1 Zero-shot retrieval effectiveness on BEIR
Table 2: Investigated prompts. The systems prompt and any text string before the prompts in this table are the same as Figure 1, thus omitted. <A> denotes the model-specific assistant special token.
ID Prompts
1 Use one word to represent the passage in a retrieval task.<A>The word is: "
2 Use one word to represent the passage.<A>The word is: "
3 Use one most important word to represent the passage in a retrieval task. Make sure your word is in lowercase.<A>The word is: "
4 Use one word to represent the passage in a retrieval task.<A>
5 Use one most important word to represent the passage in a retrieval task.<A>The word is: "
6 Use one word to represent the passage in a retrieval task. Make sure your word is in lowercase.<A>The word is: "
We present our results on BEIR in Table 1. The first observation highlights that BM25 is a very strong zero-shot retrieval method, capable of outperforming LLM2Vec, based on the Llama3-8B-Instruct LLM, across numerous datasets, achieving a higher average nDCG@10 score. This outcome implies that even with a large-size LLM, bi-directional attention enabled, additional pre-training, and SimCSE-based unsupervised contrastive training, there remains a gap in transforming a decoder-only LLM into an effective retrieval method.
@@ 5.1 Zero-shot retrieval effectiveness on BEIR
Our proposed PromptReps with Llama3-8B-Instruct LLM has lower nDCG@10 scores when only using dense or sparse retrieval. However, the hybrid system (combining dense and sparse) contributes notable retrieval effectiveness improvements, surpassing BM25 and approaching the state-of-the-art E5-PTlarge. Notably, this is achieved without any form of extra training but solely relying on prompts.
Figure 2: MRR@10 scores on MS MARCO of PromptReps with different LLMs.
@@
The scaling law of LLM Kaplan et al. (2020) also applies here. When changing Llama3-8B-Instruction to Llama3-70B-Instruction, the dense and sparse retrieval effectiveness of our PromptReps further improves, with the hybrid approach surpassing E5-PTlarge.
We further note that when interpolating dense, sparse and BM25, the average nDCG@10 achieved a remarkable score of $50.6050.60$. These results demonstrate that it is possible to build a strong retrieval system with LLMs and BM25 without the need for any unsupervised or supervised training.
5.2 Sensitivity to different prompts
In the previous experiments, we always use the prompt illustrated in Figure 1. In this section, we study how different prompts impact the retrieval effectiveness. Particularly, we design six different prompts [^10] The prompt in Figure 1 is the prompt number 6 in Table 2., listed in Table 2, and conduct experiments on TREC deep learning 2019 and 2020 datasets, and MS MARCO passage retrieval dev sub-dataset. We use Llama3-8B-Instruction as the base LLM for PromptReps. The results are listed in Table 3. We also report results of Recall@1000 and other base LLMs in Appendix A.
Table 3: Retrieval effectiveness of different prompts on TREC deep learning and MS MARCO. The ID correspond to the prompt IDs list in Table 2.
ID Methods DL2019 DL2020 MSMARCO
- BM25 49.73 48.76 18.75
- LLM2Vec - - 13.61
PrompReps Llama3-8B-Instruct (ours)
1 Dense 49.26 40.28 16.26
2 Dense 43.32 31.60 12.52
3 Dense 49.20 43.90 17.49
4 Dense 0.00 0.00 0.00
5 Dense 47.19 40.17 16.02
6 Dense 50.62 43.81 17.54
1 Sparse 41.77 44.81 20.12
2 Sparse 39.90 43.10 19.13
3 Sparse 43.50 44.87 20.42
4 Sparse 21.77 20.49 7.22
5 Sparse 42.18 44.17 19.78
6 Sparse 42.25 45.60 20.85
1 Hybrid 53.67 54.35 23.68
2 Hybrid 50.65 49.25 21.76
3 Hybrid 55.64 53.83 23.86
4 Hybrid 13.47 11.81 5.06
5 Hybrid 54.16 52.06 23.25
6 Hybrid 55.58 56.66 24.62
The results demonstrate that PromptReps can achieve a similar level of retrieval effectiveness as BM25 and surpass LLM2Vec with most of the prompts. The only prompt that does not work well is prompt #4, which does not include the phrase “The word is: “” to force the LLM to generate the representative word as the next token. This is expected because, without this phrase, the first generated token would be a general token such as “The” which is not representative of the input text.
Interestingly, our results also show that LLMs have instruction-following ability in this representation generation task. For instance, comparing prompts #1 and #2, the only difference is the phrase “in a retrieval task”, and the prompt with this phrase yields higher retrieval effectiveness across all datasets. Additionally, comparing prompts #1 and #6, the difference is the phrase “Make sure your word is in lowercase”, which matches our sparse exact matching method where we first lowercase the input text. This phrase can further improve the retrieval effectiveness. Finally, using the adjective phrase “most important” in the prompt does not significantly impact the results.
Figure 3: First-word single-representations or Multi-token single-representation.
5.3 Impact of different LLMs
In this section, we explore how different base LLMs impact PromptReps. For this study, we investigate five state-of-the-art open-sourced decoder-only LLMs, covering different model sizes and models with or without instruction tuning. We use prompt #6 for all LLMs [^11] Only the model specific conversational special tokens are changed. and report MRR@10 scores on the MS MARCO datasets. The results are illustrated in Figure 2; more detailed results including on TREC deep learning datasets are reported in Appendix A.
The results show that the hybrid retrieval effectiveness of our PromptReps consistently outperforms BM25, regardless of which LLM is used, with the only exception of Mistral-7B-Instruct. When using the Mistral-7B-Instruct LLM, the dense-only retriever performs poorly. Surprisingly, implementing PromptReps with Phi-3-mini-4k-instruct achieved much higher retrieval effectiveness than that of Mistral-7B-Instruct, despite having far less parameters (3.8B).
Meta-Llama-3 models are generally very effective for our method. For 8B models, the instruction-tuned model performs significantly better than the pretrained-only model, indicating that the instruction fine-tuning is helpful to further improve our method. The 70B instruction-tuned model achieved the best hybrid retrieval results, but the dense-only and sparse-only retrieval effectiveness are similar to the 8B instruction-tuned model. These results agree with the BEIR results presented in Table 1.
Figure 4: Multi-representations with ColBERT scoring.
Figure 5: Hybrid retrieval results of different representation methods on BEIR.
6 Alternative representations and scoring
In the previous sections, we only considered using the representations (dense and sparse) yielded from the last token in the prompt for document retrieval. These representations, in the context of generative LLMs, are responsible for predicting the first generated token. We define this setting as First-token single-representation. We have demonstrated that this simple way of generating representations is effective for document retrieval; however, these representations might be sub-optimal. For example, LLMs use sub-word tokenization algorithms such as SentencePiece Kudo and Richardson (2018). This tokenization might split a word into sub-words, meaning that the first generated token might just be a sub-word. Using the representation of the whole word might be a better representation than the first token representation. Additionally, previous works in multi-vector dense retrieval such as ColBERT Khattab and Zaharia (2020) demonstrated that using multiple representations could be beneficial for document retrieval. How can we use PromptReps to also generate single-word representations or multiple representations that can potentially enhance the retrieval effectiveness? In this section, we explore these alternative representations.
First-word single-representation and Multi-token single-representation. Instead of just using the representations for the first generated token, these two methods let the LLM finish the generation [^12] We simply use greedy generation. of the whole word or multiple words, controlled by the given prompt (“Use one word” or “Use three words”), as illustrated in Figure 3. The end of generation is detected by the token ‘’. We then pool all the representations of the generated tokens to form a single dense and sparse representation to index the input text. For the dense representation we use mean pooling and for the sparse representation we use max pooling. Once representations are obtained, the scoring is the same as First-token single-representation.
Multi-token multi-representation and Multi-word multi-representation. Instead of using a single representation for retrieval, these two methods prompt the LLM to generate multiple words and then index each generated representation separately. The difference between the two is that Multi-token multi-representation keeps all the token representations in the index, while Multi-word multi-representation first groups tokens into words by using space, and then creates a single representation for each word by using max pooling for sparse representations and mean pooling for dense representations. During retrieval, we follow the ColBERT scoring method where the relevance score of a document is computed by the sum of the maximum similarity of each query representation against each document representation (Figure 4).
Hybrid retrieval results are shown in Figure 5, and full dense and sparse retrieval results in Appendix B. Results show that all the explored methods are able to perform document retrieval. The First-token single-representation and Multi-token single-representation generally perform the best. However, we note that Multi-token single-representation requires more token generation steps and thus has higher query latency. The First-word single-representation performs the worst, suggesting that sub-word representations hurt the retrieval performance for single-word generation prompts. On the other hand, multi-representation methods with ColBERT scoring methods do not seem beneficial. Thus, we conclude that the simplest First-token single-representation is sufficient to represent the input text for document retrieval.
@@ 7 Conclusion
We introduced PromptReps, a simple yet effective method that prompts LLMs to generate dense and sparse representations for zero-shot document retrieval without any further unsupervised or supervised training. Our work reveals that modern LLMs are effective text encoders by themselves, and prompt engineering is sufficient to stimulate their text encoding ability.
For future works, techniques like few-shot in-context learning Brown et al. (2020), chain-of-thought prompting Wei et al. (2022), and auto-prompt optimization methods Yang et al. (2024); Fernando et al. (2023), which have proven to be effective in text-generation tasks, could potentially be leveraged here to enhance embedding generation.
Moreover, it has been shown that the instruction-following ability of LLMs could be transferred to embedding models with synthetic instruction fine-tuning data Wang et al. (2024b). In our work, we always keep the instruction prompt consistent across different IR tasks, which could be sub-optimal. It is interesting to investigate how to customize instructions for PromptReps to generate embeddings specific to different domains, tasks, or even to multi-lingual and cross-lingual IR settings.
Finally, our prompting method could be seen as a simple approach to obtaining a better initialization of LLM-based embedding models, which is much more cost-effective than methods requiring further pre-training BehnamGhader et al. (2024); Li et al. (2023). All the previous contrastive pre-training with paired text data and synthetically generated data could be applied on top of our method and could potentially yield improved LLM-based embedding models.
8 Limitations
PromptReps has higher query latency than other LLM-based dense retrievers if no further optimization is implemented. This limitation comes from two aspects.
First, although the computation of document representations happens offline thus will not affect query latency, the query representations are created online. PromptReps adds extra prompt texts on top of the query text thus has a longer input length – and LLM inference time is proportional to prompt length. However, we believe this limitation can be mitigated by leveraging recent works on prompt compression to compress the fixed prompt tokens into few or even a single latent token Ge et al. (2024); Cheng et al. (2024).
Secondly, the highest effectiveness for our PromptReps is achieved in the hybrid retrieval setting. Compared to previous works which use dense representations only, the hybrid setting requires both dense and sparse retrieval, thus the extra sparse retrieval introduces extra query latency (and requires additional disk/memory space for the inverted index). However, PromptReps actually only requires a limited query latency overhead if dense and sparse retrieval are implemented in parallel. In our method, obtaining both dense and sparse representations only requires a single LLM forward inference; the only extra computation is the dot product of the dense vector with the token embeddings, which is very fast on GPU. For document search, since we heavily sparsified the sparse representation, in our experiments, our sparse retriever is much faster than BM25, and the bottleneck is the dense retriever. Since the dense and sparse search could be run in parallel and the hybrid operation is a simple linear interpolation of both rankings (very fast on CPU), the query latency of the hybrid process only depends on the dense retrieval latency, and it is thus very close to previous methods.
9 Ethical considerations
In our experiments, we use PromptReps coupled with LLMs with a large number of parameters (up to 70B in our experiments) to encode the BEIR and MS MARCO datasets, which contain millions of documents. Although no LLM training was conducted, we are aware that our experiments might still have consumed significant energy, thus contributing to CO2 emissions Scells et al. (2022) and water consumption Zuccon et al. (2023).
In addition, since we leverage LLMs in a black-box manner and LLMs’ generation might contain biases Gallegos et al. (2024), the representations generated by LLMs may be biased towards certain contents or topics. Future work could consider how to mitigate biases in PromptReps via prompt engineering.
Appendix A Full results on TREC deep learning and MS MARCO
In Table 4 we present the full results we abstained on TREC deep learning datasets and MS MARCO passage retrieval dataset. The prompt ID is refer to Table 2.
Table 4: TREC deep learning and MS MARCO performance of different prompts and LLMs.
Prompt DL2019 DL2020 MS MARCO Dev
ID Methods nDCG@10 Recall@1000 nDCG@10 Recall@1000 MRR@10 Recall@1000
- BM25 49.73 74.50 48.76 80.31 18.75 85.73
- LLM2Vec - - - - 13.61 94.70
Phi-3-mini-4k-instruct (3.8B)
1 Dense 46.78 70.10 42.84 67.60 15.45 82.68
2 Dense 34.64 55.15 30.62 50.47 10.85 66.04
3 Dense 49.62 75.79 43.21 71.87 15.78 86.24
4 Dense 39.12 60.77 28.33 57.20 9.26 72.31
5 Dense 43.94 72.51 39.00 70.57 13.50 83.08
6 Dense 40.77 62.05 37.20 58.39 14.64 79.29
1 Sparse 41.51 69.56 40.95 69.70 16.89 84.72
2 Sparse 40.67 60.59 39.36 61.58 16.38 75.43
3 Sparse 42.28 74.33 40.72 72.14 18.16 87.09
4 Sparse 38.05 65.15 34.33 64.13 14.75 78.59
5 Sparse 40.68 70.84 39.26 69.02 16.04 84.41
6 Sparse 41.98 71.20 41.99 69.66 18.19 86.55
1 Hybrid 53.04 79.99 52.76 77.64 21.61 92.21
2 Hybrid 50.51 69.75 43.92 66.47 19.22 81.35
3 Hybrid 55.53 81.68 51.35 79.49 21.76 93.53
4 Hybrid 48.53 76.29 40.37 73.64 18.23 87.18
5 Hybrid 52.08 80.16 50.52 79.30 20.30 92.37
6 Hybrid 51.10 75.84 49.24 73.98 22.06 91.41
Meta-Llama-3-8B-Instruct
1 Dense 49.26 73.03 40.28 68.77 16.26 81.96
2 Dense 43.32 64.77 31.60 61.35 12.52 73.89
3 Dense 49.20 71.69 43.90 69.96 17.49 84.50
4 Dense 0.00 0.00 0.00 0.00 0.00 0.04
5 Dense 47.19 72.00 40.17 66.71 16.02 82.56
6 Dense 50.62 73.01 43.81 68.39 17.54 82.91
1 Sparse 41.77 67.28 44.81 71.36 20.12 85.71
2 Sparse 39.90 66.00 43.10 69.08 19.13 83.74
3 Sparse 43.50 66.74 44.87 72.93 20.42 85.14
4 Sparse 21.77 41.94 20.49 50.51 7.22 56.35
5 Sparse 42.18 67.18 44.17 71.94 19.78 85.37
6 Sparse 42.25 66.58 45.60 72.82 20.85 85.57
1 Hybrid 53.67 83.52 54.35 78.42 23.68 92.84
2 Hybrid 50.65 80.31 49.25 76.64 21.76 90.12
3 Hybrid 55.64 81.90 53.83 79.15 23.86 92.99
4 Hybrid 13.47 37.81 11.81 45.22 5.06 50.50
5 Hybrid 54.16 82.06 52.06 78.70 23.25 92.77
6 Hybrid 55.58 83.44 56.66 79.14 24.62 93.11
Meta-Llama-3-8B
6 Dense 43.90 67.38 35.50 63.34 14.67 79.61
6 Sparse 38.41 64.83 43.34 67.57 18.82 82.63
6 Hybrid 51.13 77.07 46.34 75.42 22.31 90.87
Mistral-7B-Instruct-v0.2
6 Dense 13.96 27.26 16.77 26.69 5.61 40.27
6 Sparse 39.84 58.05 37.29 63.53 15.62 77.55
6 Hybrid 32.58 57.00 32.95 63.12 13.18 77.98
Meta-Llama-3-70B-Instruct
6 Dense 51.95 77.30 45.01 73.66 17.76 85.65
6 Sparse 44.07 68.60 44.14 70.99 20.70 86.42
6 Hybrid 58.39 86.22 59.17 81.57 25.66 93.75
6 + BM25 63.18 88.56 62.55 86.28 27.63 95.83
Appendix B Full results of different representation methods
In Table 5 we present the full results of different representation and scoring methods discussed in Section 6.
Table 5: Full results of different representation and scoring methods on BEIR.
Dataset First token single rep First-word single rep Multi token single rep Multi-token multi-rep Multi-word multi-rep
Dense Sparse Hybrid Dense Sparse Hybrid Dense Sparse Hybrid Dense Sparse Hybrid Dense Sparse Hybrid
arguana 29.70 22.85 33.32 20.54 24.59 23.80 41.78 24.46 42.61 36.69 23.03 35.19 36.47 24.13 34.96
climatefever 19.92 9.98 21.38 13.88 11.28 16.67 22.19 9.29 20.90 19.40 6.72 17.56 18.75 8.10 18.09
dbpedia 31.53 28.84 37.71 22.71 28.70 30.08 31.83 26.33 36.03 27.46 18.18 31.35 24.50 21.66 30.32
fever 56.28 52.35 71.11 40.97 57.10 61.20 50.49 51.36 64.13 44.53 30.31 54.04 38.81 37.95 52.97
fiqa 27.11 20.33 32.40 17.61 19.60 24.74 28.94 20.73 32.44 25.26 19.41 28.38 26.28 19.50 28.30
hotpotqa 19.64 44.75 47.05 10.35 46.25 37.79 29.94 46.50 51.50 23.80 39.39 46.38 21.93 40.25 43.68
nfcorpus 29.56 28.18 32.98 21.98 29.15 29.49 28.97 28.65 33.65 25.68 25.39 31.20 22.95 25.37 30.05
nq 34.43 29.55 43.14 22.83 29.25 33.18 35.09 25.88 39.36 31.36 23.28 35.38 30.55 22.78 35.95
quora 72.55 68.27 80.45 51.68 66.68 69.09 78.90 66.83 81.93 72.71 63.21 77.91 73.57 63.38 78.77
scidocs 18.51 11.57 17.59 12.73 12.05 14.97 18.12 11.83 17.39 16.13 11.08 15.68 15.85 11.40 15.44
scifact 52.68 58.48 65.71 26.66 58.75 51.59 52.55 59.32 63.75 45.53 54.23 58.53 47.05 53.21 61.18
trec-covid 59.52 54.59 69.25 51.00 55.04 63.53 63.28 51.73 69.16 60.97 49.88 63.54 61.17 46.24 65.10
touche2020 14.85 18.47 21.65 12.23 21.58 19.13 15.59 19.24 21.44 15.86 17.81 22.10 15.41 18.09 19.26
avg 35.87 34.48 44.13 25.01 35.39 36.56 38.28 34.01 44.18 34.26 29.38 39.79 33.33 30.16 39.54